OpenStudy (anonymous):

Suppose the population of a town is 16,400 and is growing 4% each year. a. Write an equation to model the population growth. b. Predict the population after 12 years.

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

I believe par a would be \[16,400x.04^{x}\], x standing for the amount of years that pass so for part b plug 12 in for x. If you still dont get it just let me know

OpenStudy (anonymous):

so what is it

OpenStudy (wolf1728):

Population = 16,400*(1.04)^years Population = 16,400*(1.04)^12 Population = 26,257

OpenStudy (anonymous):

Yupp

OpenStudy (anonymous):

waht bout a part

OpenStudy (wolf1728):

Population = 16,400*(1.04)^years

OpenStudy (anonymous):

Population=16,400x(1.04)^number of years

OpenStudy (wolf1728):

It was the first line I typed

OpenStudy (wolf1728):

Gee thanks wolf Shucks you are welcome

OpenStudy (anonymous):

Lol ya did good wolf..you got a new fan here

OpenStudy (wolf1728):

Thanks Ihatescholl - although I would have appreciated it if the person who asked the question were as thankful :-)